# BPE: Behavioral Profiling Ensemble

### v1

@@ 4.1.2 Base Model Pool
We assembled a diverse repository containing 13 heterogeneous base classifiers, spanning the mainstream algorithmic paradigms in machine learning:
@@
    Non-linear and Distance-based Models: Including MLPClassifier [34], SVC (with probability estimation) [22], and K-Neighbors (KNN) [23].
@@
To ensure experimental validity, we implemented a strict performance-based pruning mechanism. We retain only those base models $h_{k}$ whose error rate $E(h_{k})$ satisfies:
$E(h_{k})\leq E_{\text{best}}\times(1+\alpha)$ (10)
where $E_{\text{best}}$ denotes the error of the best-performing model on the validation set, and the tolerance threshold is set to $\alpha=0.15$. This mechanism effectively filters out suboptimal models (“weak links”) that are ill-suited for the current data distribution[36].
It should be noted that this setup is only for the purpose of the automated base model pool, because in actual deployment we do not include models with too large a gap in the integration. The validation set here is only used to simulate this situation.
@@ 4.1.3 Baseline Methods
Static Ensemble Baselines.
@@
    Simple Average: This method operates on the assumption that all base models contribute equally. It computes the arithmetic mean of the output probability vectors from all base classifiers as the final prediction. Despite its simplicity, consistent with Dietterich’s statistical perspective, simple averaging effectively smooths decision boundaries and reduces variance by aggregating multiple independent hypotheses, serving as a robust baseline that is often challenging to outperform significantly.
@@
    Weighted Average: An empirical strategy derived from engineering practice. In industrial applications and data science competitions, the Out-of-Fold (OOF) accuracy generated during the cross-validation phase is typically utilized as static weights. This method relies on the intuitive assumption that models demonstrating superior performance on the validation set should be granted higher authority during the testing phase.
@@
    Rank Average: A robust aggregation technique widely adopted to address probability calibration discrepancies in heterogeneous ensembles. To mitigate the interference caused by scale inconsistencies across different algorithms, Rank Average discards absolute probability values in favor of fusing the relative rankings of predicted scores. It is widely regarded as a highly stable engineering baseline.
@@
    Median Average: In scenarios where output probability vectors exhibit extreme outliers, the median offers superior robustness compared to the mean and serves as a commonly utilized ensemble technique.
@@
    Stacking: A meta-learning-based ensemble framework proposed by Wolpert [5]. In this study, logistic regression is employed as the meta-learner. The training process utilizes the OOF predictions of base models on the validation set as meta-features and ground-truth labels as targets to learn a set of linear combination weights that minimize global loss. While Stacking can automatically capture the complementarity between models, its learned weights remain fixed during inference and cannot adapt dynamically to individual test samples.

### v2

@@ 5.2 Base learner Pool
We assembled a diverse pool containing 13 heterogeneous base classifiers, spanning the mainstream algorithmic paradigms in machine learning:
@@
    Non-linear and Distance-based Models: Including MLPClassifier [37], SVC (with probability estimation) [24], and k-Nearest Neighbors (KNN) [25].
@@
In practical applications, an ensemble system typically does not use exactly the same model pool for all datasets. Prior work has also pointed out that overly weak models may harm ensemble performance [39, 40, 41, 42]. Consistent with this insight, our experiments show that when no screening is performed and obviously mismatched models are forcibly included, the classification accuracy of all evaluated algorithms decreases. Therefore, it is common in practice to construct an appropriate candidate model pool by combining prior knowledge with a screening strategy; for example, in classical machine learning, candidate algorithms are selected according to whether the data satisfy certain model assumptions, and in data science competitions, models with clearly low accuracy are often excluded based on public leaderboards or cross-validation performance.
To simulate this practical workflow, we introduce a simple performance-based screening mechanism as a unified experimental protocol for model-pool construction. Concretely, for each dataset we (i) split the training set into 80%/20% for screening, (ii) compute the screening accuracy $\mathrm{ACC}(h_{k})$ of each base learner on the 20% split, (iii) retain only the models satisfying $\mathrm{ACC}(h_{k})\geq\mathrm{ACC}_{\text{best}}\times(1-\alpha)$ with $\alpha=0.15$, and (iv) retrain the retained models on the full training set for subsequent evaluation. Besides mimicking real-world workflows, this dataset-dependent base model pool introduces diversity across datasets, which makes the empirical comparison more robust. Meanwhile, for each dataset, all ensemble methods are evaluated using the same screened model pool, ensuring a fair comparison between algorithms. We emphasize that this screening mechanism is part of the experimental protocol for model-pool construction; the BPE framework itself does not rely on a validation set for instance-wise weighting.
@@ 5.3 Baseline Methods
Our experiments involve multiple conventional DES/DCS baselines that require a reference set. We observe that splitting an additional fixed reference set from the training data significantly reduces the data available for training base models, which in turn leads to a clear drop in classification accuracy for these reference-dependent methods. To ensure these methods can utilize the full dataset and achieve accuracy that is comparable with other methods, we adopt their most common standard implementation in practice: using Out-of-Fold (OOF) predictions to fully exploit all training data [7]. Concretely, we perform $5$-fold OOF, where base models are trained on each fold and used to generate OOF predictions on the held-out fold; after OOF predictions are obtained, each base model is retrained on the full training set. Note that $5$-fold OOF introduces additional training cost, and therefore creates a significant computational gap between these reference-dependent methods and BPE.
In addition, to present the impact of different protocols, we report in the Appendix the results of a standard engineering alternative: “fixed train/reference split + retraining on the full training set after training.” This alternative typically costs only about $2\times$ the computation of BPE, which is smaller than $5$-fold OOF, but its accuracy is generally worse than the OOF-based strategy.
Static Ensemble Baselines
@@
    Single Best Classifier: (abbreviated as SB in the following tables) This oracle-style reference baseline selects the single best-performing base model according to the reference accuracy used in the screening protocol, and then evaluates this selected model on the test set.
@@
    Simple Average: (abbreviated as SA in the following tables) This method operates on the assumption that all base models contribute equally. It computes the arithmetic mean of the output probability vectors from all base classifiers as the final prediction. Despite its simplicity, consistent with Dietterich’s statistical perspective, simple averaging effectively smooths decision boundaries and reduces variance by aggregating multiple independent hypotheses, serving as a robust baseline that is often difficult to outperform significantly [43, 2].
@@
    Weighted Average: (abbreviated as WA in the following tables) An empirical strategy derived from engineering practice. In industrial applications and data science competitions, the Out-of-Fold (OOF) accuracy generated during the cross-validation phase is typically utilized as static weights. This method relies on the intuitive assumption that models demonstrating superior performance on the reference set should be assigned higher weights. during the testing phase.
@@
    Median Average: (abbreviated as MA in the following tables) In scenarios where output probability vectors exhibit extreme outliers, the median offers superior robustness compared to the mean and serves as a commonly utilized ensemble technique.
Dynamic Ensemble Baselines
  • 1.
  • 2.
  • 3.
  • 4.
@@
  • 6.
  • 7.
    DES-KNN: This family of methods selects an ensemble for each test instance by jointly considering local accuracy and diversity within its Region of Competence (RoC) [45]. Concretely, base classifiers are ranked by decreasing accuracy and (separately) by increasing redundancy/diversity estimated on the RoC. The final Ensemble of Competence (EoC) is formed by combining the top-$p_{a}$ most accurate models with the top-$p_{b}$ most diverse models (with duplicates removed). In our implementation, the diversity score of a classifier is obtained from its average pairwise diversity to the remaining classifiers on the RoC, using common measures such as Double-Fault (DF), the $Q$-statistic, and the Ratio of Errors (RE).